Skip to Search

diagonal spacer

Systems Staff Meeting Notes
April 17, 2006

Present: Cynthia Bowker, Will Chaney, Suzy Covey, Debra Fetzer, Tom Minton, Chris Nicolich, Geoffrey Stewart, Ying Tang, Stephen Williams
  1. Rich is out this week. Since Winston is out at training all week, Suzy will be the ALEPH back-up person.
  2. Rob Roberts - got the second quote for the carpet replacement. Set a tentative date of August 1 to begin abatement. Need to wait until folks are back in West.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. 15 quotes from Dell are about ready to go. We should be able to start ordering today. Dell cut us a good deal; we should be under budget, depending on the software costs. Software order to follow hardware.
  2. The OmniView monitors should ship out by the end of this week.
  3. The UPS battery for room 16 - we just got a bunch of boxes from CDWG - it could be down there right now. Need to change it out ASAP.
  4. Geoffrey delivered a 17" monitor to West so they could see how it fits on the mounting brackets and shelf for the dual view, flat-paneled monitors for the Circulation desk on the second floor of West. It needs to be modified, because you could not get the monitor in or out.
  5. Cynthia notes that within the last two months, there have been for motherboards in need of replacement. They failed due to bulging capacitors for the second time. All were still under warranty. Dell GX270's.
  1. Exchange outage - three services: Exchange Domain Service, Enterprise services & Services - theses three security groups have to remain in the user container. The server had to reboot on Thursday after updates ran. In doing so, the server redid the registry. An error code in the log hinted to this - Will called Dell and the event lead us to finding out the fix.
  2. Campus AD - summer.
Software Projects Status
  1. Things are the same with the ERES interface - Chris hopes to finish this week or next. They want to pilot it in May. However, the lead contact is out on maternity leave.
  2. NT4 Indexing algorithm - Ying says it is done. Need to do some more work to clear up duplicates. Mark says they will probably want to get rid of the duplicates. Ultimately, Erich will decide.
  3. Grant Tracking database - Stephen says it is almost done. Want to have Winston check on this before he leaves for vacation next week.
  4. Patron-update can be removed from the list.
  5. Greenstone loader is now working. It runs once a day. Pre-loader runs twice a day.
Web Projects
  1. Circ manual to be protected by IP address. Need to get Health & Law IPs. Carol Turner wants all staff and public machines to be able to access.
  2. Metalib meeting next Monday. We are going to go with it - for a summer soft launch.
  3. Homepage to include photos of the new library for the opening of West.
DP Coordinator
  1. Students - last day of spring semester classes is Wed April 26. Finals are the first week of May. There's a week break before the start of Summer A on May 15.
  2. All three students will be returning for the summer. Schedule is in progress.


University of Florida Home Page